OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 –3x + 2y = 1 2 4 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (zepp):

Reduction method, add them up so 3x could be eliminated y+2y=11+1 3y=12 Now solve you y :)

OpenStudy (anonymous):

thanks that helped alot

OpenStudy (zepp):

You are welcome :)

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system. Question 5 (Multiple Choice Worth 2 points) What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3 −1 5 −2 2

OpenStudy (zepp):

Post other questions on the left, so other people could earn medals as well. :)

OpenStudy (anonymous):

ok